Did I do this composite function correctly?
Let f(x) = 1/(2x), g(x) = 7x^3, and h(x) = −4x^2 + 1
(f º g º h)(x) = 1/(686(-4x^2+1)^3)
Is that right?
Thank You

Answer by Theo(13342)

i don't believe so.

i got something entirely different and i confirmed it is correct.

start with:

f(x) = 1/(2x)
g(x) = 7x^3
h(x) = -4x^2 + 1

(f.g.h)(x) = f(g(h(x)))

that translates to f of g of h of x.

you find h(x) first.
then you find g(h(x)) next.
then you find f(g(h(x))) next.

h(x) = (-4x^2 + 1)

g(h(x)) = g(-4x^2 + 1) = (7 * (-4x^2 + 1)^3)

f(g(h(x)) = f(7 * (-4x^2 + 1)^3) = 1 / (2 * (7 * (-4x^2 + 1)^3))

if you work it all through, you will get:

f(g(h(x)) = 1 / (-896x^6 + 672x^4 - 168x^2 + 2)

the piece parts are:

h(x) = -4x^2 + 1

g(h(x)) = -448x^6 + 336x^4 - 84x^2 + 1

f(g(h(x)) = 1 / (-896x^6 + 672x^4 - 168x^2 + 2)

i confirmed by assigning the arbitrary value of 7 to x and then solving from the original equations and then solving from the final equation.

i got:

h(7) = -195

g(-195) = -51904125

f(-51904125) = -9.63314573*10^-9

I then went to the final equation to get:

(f.g.h)(7) = f(g(h(7)) = -9.63314573*10^-9

since both answers are the same, i assume that i did it right.

you can find (-4x^2 + 1)^3 by first multiplying (-4x^2 + 1) by (-4x^2 + 1) and the multiplying the result by (-4x^2 + 1).

you can also find it by seeing that it is a binomial expansion of (a + b)^3 where a = -4x^2 and b = 1.

i did it both ways and got the same answer of (-4x^2 + 1)^3 = -64x^6 + 48x^4 - 12x^2 + 1